OpenStudy (anonymous):

Which set of numbers could represent the lengths of the sides of a right triangle? Answers: 4, √35 , 9 6, √82 , 13 5, √55 , 10 11, √47, 13

terenzreignz (terenzreignz):

Here... a systematic way to do it is to take a and b, the two smaller numbers, and consider this interval : \[\Large \left(|a-b| \ , \ a+b\right)\] And if the third side falls within that interval, then it's a legit triangle...

OpenStudy (anonymous):

@terenzreignz Right triangle.

terenzreignz (terenzreignz):

Pardon?

terenzreignz (terenzreignz):

Oh crud... :/ Sorry, @Virgil @Aphex didn't read through...

OpenStudy (anonymous):

so whats the answer?? @terenzreignz

terenzreignz (terenzreignz):

Yeah... uhh, to figure out which represents the right triangle, take the two smaller sides, a and b, and square them \[\Large a^2 + b^2\] The result should be the square of the the third side for it to be a right triangle...

terenzreignz (terenzreignz):

Try the first one... \[\large 4^2=16\]\[\large (\sqrt{35})^2 = 35\]\[\Large 16+35 = 51 \ne 81=9^2\] So, the first choice isn't it :)

OpenStudy (anonymous):

ok, its not the first one, im gonna take a lucky guess and say C